If the measure of the vertex angle of an isosceles triangle is 24degrees less than the sum of the measures of its base angles, then the measure of one of its base angles is:
(A) 156 degrees (B) 66 degrees (C) 24 degrees (D) 78 degrees (E) 51 degrees

Answer by mananth(16949)

If the measure of the vertex angle of an isosceles triangle is 24degrees less than the sum of the measures of its base angles, then the measure of one of its base angles is:
(A) 156 degrees (B) 66 degrees (C) 24 degrees (D) 78 degrees (E) 51 degrees
let the vertex angle be x
the sum of the other two angles = 180-x
x+24 = 180-x
2x=180-24
2x=156
x= 78
180-78 will be the aum of the other two equal angles
180-78 /2 will be the angle =51